What is the scheduled principal payment (SP) for Month 2 of a residential mortgage with an...

  1. What is the scheduled principal payment (SP) for Month 2 of a residential mortgage with an initial balance of $200,000, coupon rate =7.2%, and an original maturity of 15 years?
Answer #1

First we compute the monthly payment as below

Using financial calculator
Input:

PV = 200000

I/Y = 7.2/12

N = 15*12 = 180

Solve for PMT =1820.09

The month 1 interest = 200000*7.2%/12 = 1200

Hence year 1 principal balance = 200000- (1820.09-1200) = 199379.91

month 2 interest = 199379.91*7.2%/12 = 1196.28

Month 2 principal = 1820.09-1196.28 = $623.81

Answer = $623.81